OpenStudy (anonymous):

if Z=0.064 find the value of (1/z)^1/3

OpenStudy (anonymous):

\[\frac{ 1 }{ z }=\frac{ 1 }{ 0.064 }=\frac{ 125 }{ 8 }=\frac{ 5^3 }{ 2^3 }\] Therefore, \[\left( \frac{ 1 }{ z } \right)^{\frac{ 1 }{ 3 }}=\left( \frac{ 5^{3} }{ 2^{3} } \right)^{\frac{ 1 }{ 3 }}=\frac{ 5^{3\times \frac{ 1 }{ 3 }} }{ 2^{3\times \frac{ 1 }{ 3 }} }=\frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

You can do this without my help @tinybookworm

OpenStudy (anonymous):

You can do this without my help @tinybookworm

OpenStudy (anonymous):

@prettybunny123456 Thank you. I am sure I will need your help.

OpenStudy (anonymous):

Nah! "Believe in yourself. I believe in you."

OpenStudy (anonymous):

Sound familiar

OpenStudy (anonymous):

Good night y'all